OpenStudy (anonymous):

I really need help with algebra 1!! Systolic blood pressure, the higher number in a blood pressure reading, is measured as your heart muscle contracts. The inequality P ≤ ½a + 110 gives the normal systolic blood pressure, P, based on age, a, (in years). Find the systolic blood pressure for a 16 year old. At age , 30 does 125 represent a maximum or minimum systolic pressure? Explain:

OpenStudy (anonymous):

uuufff boy this is way too hard!

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

sorry bro

OpenStudy (anonymous):

im a chick lol but thanks anyways

OpenStudy (anonymous):

huh? a chick watcha mean?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Have you tried substituting those specific ages into your inequality?

OpenStudy (anonymous):

no, i dont know how :/

OpenStudy (anonymous):

Well, the information tells you that a stands for the age, then in part A, they give you an age of 16 so plug 16 in for a \[P \le \frac{ 1 }{ 2 }a+110 \rightarrow P \le \frac{ 1 }{ 2 }16+110\]

OpenStudy (anonymous):

oh! okay im going to try that

OpenStudy (anonymous):

its not working for me, how do you do it?

OpenStudy (anonymous):

the chossen she is a girl not a guy thats what she means and this is a problem in Algebra 1???? are you sure? its soooooooooooooooooooooooo hard

OpenStudy (anonymous):

Could you show your work?

OpenStudy (karatechopper):

Uh. Plug 16 in for a. Multiply and simplify.

OpenStudy (karatechopper):

That part it very simple.